# Long and Short Guidance in Score identity Distillation for One-Step Text-to-Image Generation

### v3

@@ 4 Conclusion
In this paper, we introduce a novel approach that combines Classifier-Free Guidance (CFG) with Score identity Distillation (SiD) to efficiently distill Stable Diffusion models into effective one-step generators. Utilizing our innovative Long and Short CFG strategies (LSG), we have distilled these models using only synthetic images generated by the one-step generator. This method not only validates the practical potential of SiD but also establishes new benchmarks for one-step diffusion distillation, achieving remarkable zero-shot FID scores on the COCO-2014 validation set. Designed for enhanced efficiency without sacrificing performance, our approach allows for learning from the teacher model without requiring real images or incorporating additional regression or adversarial losses. We have made our code and distilled models publicly available to facilitate further research.
Acknowledgments and Disclosure of Funding
We are grateful for the contributions of Michael (Qijia) Zhou in preparing the data and code. We also extend our thanks to Tianwei Yin for clarifying our queries regarding the evaluation protocol of DMD on FID and CLIP scores. Additionally, we appreciate Thuan Hoang Nguyen for responding to our questions about SwiftBrush and for kindly providing a model checkpoint upon request. M. Zhou, H. Zheng, and Z. Wang acknowledge the support of NSF-IIS 2212418 and NIH-R37 CA271186.
@@ Appendix B Related work
Generative modeling of high-dimensional data has long been a focal point in machine learning research. This area primarily concentrates on replicating various data distributions: the original data distribution, conditional distributions influenced by labels, noisy or incomplete measurements, textual descriptions, or the joint distribution of data and other modalities. This has spurred the development of a diverse range of generative models and methodologies. Initially, these models were only capable of handling simpler, low-dimensional data such as $28\times 28$ grayscale or binarized MNIST digits [Hinton et al., 2006, Salakhutdinov and Hinton, 2009, Vincent et al., 2010], vector-quantized local descriptors [Fei-Fei and Perona, 2005, Chong et al., 2009], or patches of natural and hyperspectral images [Zhou et al., 2009, Xing et al., 2012, Polatkan et al., 2015]. Early models often utilized neural networks with stochastic binary hidden layers or shallow hierarchical Bayesian models, which are simpler to train but have limited capacities.
@@
Deep generative models. To tackle the generation of high-dimensional data, such as images comprising millions of pixels, substantial advancements in generative models have been made over the past decade. This period has marked the emergence of diverse deep generative models, including variational auto-encoders (VAEs) [Kingma and Welling, 2014, Rezende et al., 2014], normalizing flows [Papamakarios et al., 2019], generative adversarial networks (GANs) [Goodfellow et al., 2014, Reed et al., 2016, Karras et al., 2019, Wang et al., 2023b], autoregressive models [Gregor et al., 2015, Mansimov et al., 2015], and diffusion models [Sohl-Dickstein et al., 2015, Song and Ermon, 2019, Ho et al., 2020, Song and Ermon, 2020, Song et al., 2021b, a, Dhariwal and Nichol, 2021, Karras et al., 2022, Peebles and Xie, 2023, Zheng et al., 2024]. Additionally, essential resources for creating effective text-to-image synthesis systems, such as large language models [Vaswani et al., 2017, Devlin et al., 2018, Radford et al., 2018, 2019, Raffel et al., 2020, He et al., 2020, Brown et al., 2020, Achiam et al., 2023, Touvron et al., 2023, Jiang et al., 2023], large vision-language models [Radford et al., 2021, Ilharco et al., 2021, Cherti et al., 2023], advanced visual tokenization and compression [Rolfe, 2016, van den Oord et al., 2017, Esser et al., 2021, Rombach et al., 2022, Podell et al., 2024], and extensive training datasets [Thomee et al., 2016, Changpinyo et al., 2021, Schuhmann et al., 2022], have transitioned from proprietary tools of major tech companies to publicly available assets.
@@
Previously, text-to-image models primarily utilized GANs and focused on small-scale, object-centric domains like flowers, birds, and face images [Reed et al., 2016, Zhang et al., 2017, Xu et al., 2018, Zhang et al., 2020]. Powered by text encoders that leverage pretrained large language models or vision-language models, which offer profound language comprehension and extract semantically rich latent representations, and supported by an extensive collection of text-image pairs, three main families of generative models—GANs [Zhou et al., 2022, Sauer et al., 2022, Kang et al., 2023], autoregressive models [Ramesh et al., 2021, Zhang et al., 2021, Ding et al., 2021, Gafni et al., 2022, Yu et al., 2022, Chang et al., 2023], and diffusion models [Nichol et al., 2022, Ramesh et al., 2022, Saharia et al., 2022, Rombach et al., 2022, Xu et al., 2022, Wang et al., 2023a, Qin et al., 2023]—have effectively capitalized on these technological advancements. These developments have facilitated the creation of text-to-image synthesis systems that demonstrate exceptional photorealism and sophisticated language understanding.
@@
Acceleration of diffusion-based generation. Early-stage diffusion models were notably slow in sampling, spurring extensive research aimed at speeding up the reverse diffusion process. Researchers have approached this by interpreting diffusion models through stochastic or ordinary differential equations and using advanced numerical solvers to enhance efficiency [Song et al., 2021c, a, Liu et al., 2022a, Lu et al., 2022b, Zhang and Chen, 2023, Karras et al., 2022]. Additional strategies include truncating the diffusion chain to initiate generation from more structured distributions [Pandey et al., 2022, Zheng et al., 2023b, Lyu et al., 2022], integrating these models with GANs to boost generation speed [Xiao et al., 2022, Wang et al., 2023b], and exploring flow matching in diffusion modeling [Liu et al., 2022b, Lipman et al., 2022, Albergo et al., 2023]. More recently, research has pivoted towards distilling reverse diffusion chains to refine and expedite the generation process, a direction that continues to evolve with new methodologies and insights [Luhman and Luhman, 2021, Salimans and Ho, 2022, Zheng et al., 2023a, Meng et al., 2023, Song et al., 2023, Song and Dhariwal, 2023, Kim et al., 2023, Sauer et al., 2023b, Xu et al., 2023, Yin et al., 2023, Luo et al., 2023c, Zhou et al., 2024].
@@
Text-to-image diffusion distillation. Recent efforts aim to accelerate the sampling process from pre-trained diffusion teachers like SD [Rombach et al., 2022]. Sauer et al. [2023b] focused on distilling diffusion models into generators capable of one or two-step operations through adversarial training. Xu et al. [2023] introduced UFOGen, utilizing a time-step-dependent discriminator for generator initialization. Luo et al. [2023a] applied consistency distillation [Song et al., 2023] to text-guided latent diffusion models [Ramesh et al., 2022] for efficient, high-fidelity text-to-image generation. Building on the idea of using a pre-trained 2D text-to-image diffusion model for text-to-3D synthesis [Poole et al., 2022, Wang et al., 2023c], SwiftBrush [Nguyen and Tran, 2024] showcases its effectiveness of distilling pre-trained stable diffusion models. Distribution Matching Distillation (DMD) by Yin et al. [2023] further enhances distillation quality by adding a regression loss term.
@@
How does SiD-LSG differ from previous methods? SiD-LSG introduces several unique features that distinguish it from previous text-to-image diffusion distillation methods. Firstly, similar to SwiftBrush [Nguyen and Tran, 2024], SiD-LSG is a data-free distillation method, which eliminates the need for original training datasets or synthetic data generated with SDE/ODE solvers during the distillation process. However, unlike SwiftBrush, SiD-LSG necessitates gradient backpropagation through the score networks, a critical step akin to gradient backpropagation through the discriminator in methods employing adversarial losses. Additionally, SiD-LSG applies CFG to both the training and evaluation of the fake score network. This is a departure from previous methods, which typically apply CFG only during the evaluation of the pretrained score network. Finally, SiD-LSG aims to minimize a model-based explicit score-matching loss, a type of Fisher divergence. In contrast, previous methods often focus on minimizing losses based on KL divergence, consistency, regression, GAN-based adversarial tactics, or a combination thereof.
Appendix C Training and evaluation details
@@
The hyperparameters tailored for our study are outlined in Table 3. It’s important to note that the time and memory costs reported in Table 3 do not account for those incurred during periodic evaluations of the FID and CLIP scores of the single-step generator, nor do they include the resources used to save model checkpoints during the distillation process. These costs can vary significantly depending on the computing and storage platforms used and the frequency of these operations.
@@
Appendix D Limitations and computational requirements
Memory and speed. Table 3 in Appendix C offers a detailed examination of the computational resources required for the SiD distillation process employing various CFG strategies—Long, Short, basic Long-Short, and the recommended Long-Short—across different NVIDIA GPU platforms (RTX-A5000 with 24GB, RTX-A6000 with 48GB, and H100 with 80GB). Key observations include:
SiD-LSG can be effectively run on the RTX-A5000 with 24GB memory by enabling xFormers [Lefaudeux et al., 2022] and switching to FP16 for model and gradient precision. On the RTX-A6000 with 48GB memory, it operates under FP32 with xFormers enabled. However, on our available H100 with 80GB memory, where xFormers are not supported, there is a noticeable increase in GPU memory consumption.
@@ Appendix D Limitations and computational requirements
Further investigation is required to optimize FP16 performance to equal that of FP32, potentially through refined loss scaling techniques or by implementing an effective warmup period using FP16 before transitioning to FP32. We plan to address these challenges in future studies.
Reaching a performance plateau. Zhou et al. [2024] demonstrate through extensive comparisons that the SiD distilled one-step generator can reduce the FID at an exponential decay rate, showing a log-log linear relationship between the number of iterations and FID. This approach can match or even surpass the performance of both unconditional and label-conditional teacher diffusion models trained under the EDM framework [Karras et al., 2022], provided sufficient training.
However, results shown in Figures 4 and 5 indicate that the SiD-LSG distilled one-step generator on SD1.5 quickly reaches a performance plateau in reducing FID and/or increasing CLIP scores, particularly when the LSG scale exceeds 2. Additionally, data from Table 1 demonstrate that after processing 10M images (approximately 20k iterations at a batch size of 512), SiD-LSG still does not match the text-image alignment performance of the teacher model, which achieves higher CLIP scores after 250 generation steps. Notably, by reducing the LSG scale to 1.5 and doubling the training duration, SiD-LSG achieves a record-low FID of 8.15, establishing a new benchmark among all diffusion distillation methods. This achievement also surpasses the teacher model’s FID of 8.78, which was obtained with a CFG scale of 3 and 250 generation steps. However, this reduction in the LSG scale to 1.5 also results in a significant decline in its CLIP score.
This suggests significant potential for further performance enhancements, possibly by extending beyond single-step generation, increasing the model size, or incorporating real data and additional regression or adversarial losses into the distillation process. These avenues for improvement will be the focus of future studies.
Appendix E Prompts and additional example images

### v4

@@ 4 Conclusion and Future Work
This paper introduces a novel data-free method combining Classifier-Free Guidance (CFG) with Score identity Distillation (SiD) to efficiently distill Stable Diffusion models into effective one-step generators. By leveraging our innovative Long and Short CFG strategies (LSG), we distilled these models using only synthetic images generated by the one-step generator. This approach not only validates the practical potential of SiD but also sets new benchmarks for data-free one-step diffusion distillation, achieving remarkable zero-shot FID scores on the COCO-2014 validation set. Our method enhances efficiency while maintaining generation performance, allowing learning from the teacher model without the need for real images or the inclusion of additional regression or adversarial losses. We will make our code and distilled models publicly available to facilitate further research.
In the data-free setting, we are exploring the use of SiD-LSG for privacy and security-sensitive tasks where access to actual training data is not feasible. While we have advanced the capabilities of data-free diffusion distillation, our baseline methods such as ADD and DMD, typically require the use of real or teacher-synthesized images. Moving forward, we plan to lift the data-free constraint and integrate SiD-LSG with Diffusion GAN-based adversarial training, which has successfully transformed pretrained unconditional and label-conditional diffusion models into one-step generators, achieving state-of-the-art generation performance without CFG (Zhou et al., 2025). This transition will involve the use of real images to not only further enhance photo-realism and improve text alignment but also adapt SiD-LSG distilled one-step generators to domains that differ from those used to train the teacher. This approach aims to broaden the applicability and effectiveness of our SiD-LSG distilled diffusion models.
Acknowledgments
M. Zhou, Z. Wang, and H. Zheng acknowledge the support of NSF-IIS 2212418 and NIH-R37 CA271186.
@@ Appendix B Related work
Generative modeling of high-dimensional data has long been a focal point in machine learning research. This area primarily concentrates on replicating various data distributions: the original data distribution, conditional distributions influenced by labels, noisy or incomplete measurements, textual descriptions, or the joint distribution of data and other modalities. This has spurred the development of a diverse range of generative models and methodologies. Initially, these models were only capable of handling simpler, low-dimensional data such as $28\times 28$ grayscale or binarized MNIST digits (Hinton et al., 2006; Salakhutdinov & Hinton, 2009; Vincent et al., 2010), vector-quantized local descriptors (Fei-Fei & Perona, 2005; Chong et al., 2009), or patches of natural and hyperspectral images (Zhou et al., 2009; Xing et al., 2012; Polatkan et al., 2015). Early models often utilized neural networks with stochastic binary hidden layers or shallow hierarchical Bayesian models, which are simpler to train but have limited capacities.
@@
Deep generative models. To tackle the generation of high-dimensional data, such as images comprising millions of pixels, substantial advancements in generative models have been made over the past decade. This period has marked the emergence of diverse deep generative models, including variational auto-encoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014), normalizing flows (Papamakarios et al., 2019), generative adversarial networks (GANs) (Goodfellow et al., 2014; Reed et al., 2016; Karras et al., 2019; Wang et al., 2023b), autoregressive models (Gregor et al., 2015; Mansimov et al., 2015), and diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Song & Ermon, 2020; Song et al., 2021b; a; Dhariwal & Nichol, 2021; Karras et al., 2022; Peebles & Xie, 2023; Zheng et al., 2024). Additionally, essential resources for creating effective T2I synthesis systems, such as large language models (Vaswani et al., 2017; Devlin et al., 2018; Radford et al., 2018; 2019; Raffel et al., 2020; He et al., 2020; Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2023), large vision-language models (Radford et al., 2021; Ilharco et al., 2021; Cherti et al., 2023), advanced visual tokenization and compression (Rolfe, 2016; van den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022; Podell et al., 2024), and extensive training datasets (Thomee et al., 2016; Changpinyo et al., 2021; Schuhmann et al., 2022), have transitioned from proprietary tools of major tech companies to publicly available assets.
@@
Previously, T2I models primarily utilized GANs and focused on small-scale, object-centric domains like flowers, birds, and face images (Reed et al., 2016; Zhang et al., 2017; Xu et al., 2018; Zhang et al., 2020). Powered by text encoders that leverage pretrained large language models or vision-language models, which offer profound language comprehension and extract semantically rich latent representations, and supported by an extensive collection of text-image pairs, three main families of generative models—GANs (Zhou et al., 2022; Sauer et al., 2022; Kang et al., 2023), autoregressive models (Ramesh et al., 2021; Zhang et al., 2021; Ding et al., 2021; Gafni et al., 2022; Yu et al., 2022; Chang et al., 2023), and diffusion models (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Xu et al., 2022; Wang et al., 2023a; Qin et al., 2023)—have effectively capitalized on these technological advancements. These developments have facilitated the creation of T2I synthesis systems that demonstrate exceptional photorealism and sophisticated language understanding.
@@
Acceleration of diffusion-based generation. Early-stage diffusion models were notably slow in sampling, spurring extensive research aimed at speeding up the reverse diffusion process. Researchers have approached this by interpreting diffusion models through stochastic or ordinary differential equations and using advanced numerical solvers to enhance efficiency (Song et al., 2021c; a; Liu et al., 2022a; Lu et al., 2022b; Zhang & Chen, 2023; Karras et al., 2022). Additional strategies include truncating the diffusion chain to initiate generation from more structured distributions (Pandey et al., 2022; Zheng et al., 2023b; Lyu et al., 2022), integrating these models with GANs to boost generation speed (Xiao et al., 2022; Wang et al., 2023b), and exploring flow matching in diffusion modeling (Liu et al., 2022b; Lipman et al., 2022; Albergo et al., 2023). More recently, research has pivoted towards distilling reverse diffusion chains to refine and expedite the generation process, a direction that continues to evolve with new methodologies and insights (Luhman & Luhman, 2021; Salimans & Ho, 2022; Zheng et al., 2023a; Meng et al., 2023; Song et al., 2023; Song & Dhariwal, 2023; Kim et al., 2023; Sauer et al., 2023b; Xu et al., 2023; Yin et al., 2023; Luo et al., 2023c; Zhou et al., 2024).
@@
T2I diffusion distillation. Recent efforts aim to accelerate the sampling process from pre-trained diffusion teachers like SD (Rombach et al., 2022). Sauer et al. (2023b) focused on distilling diffusion models into generators capable of one or two-step operations through adversarial training. Xu et al. (2023) introduced UFOGen, utilizing a time-step-dependent discriminator for generator initialization. Luo et al. (2023a) applied consistency distillation (Song et al., 2023) to text-guided latent diffusion models (Ramesh et al., 2022) for efficient, high-fidelity T2I generation. Building on the idea of using a pre-trained 2D T2I diffusion model for text-to-3D synthesis (Poole et al., 2022; Wang et al., 2023c), SwiftBrush (Nguyen & Tran, 2024) showcases its effectiveness of distilling pre-trained stable diffusion models. Distribution Matching Distillation (DMD) by Yin et al. (2023) further enhances distillation quality by adding a regression loss term.
@@
How does SiD-LSG differ from previous methods? SiD-LSG introduces several unique features that distinguish it from previous T2I diffusion distillation methods. Firstly, similar to SwiftBrush (Nguyen & Tran, 2024), SiD-LSG is a data-free distillation method, which eliminates the need for original training datasets or synthetic data generated with SDE/ODE solvers during the distillation process. However, unlike SwiftBrush, SiD-LSG necessitates gradient backpropagation through the score networks, a critical step akin to gradient backpropagation through the discriminator in methods employing adversarial losses. Additionally, SiD-LSG applies CFG to both the training and evaluation of the fake score network. This is a departure from previous methods, which typically apply CFG only during the evaluation of the pretrained score network. Finally, SiD-LSG aims to minimize a model-based explicit score-matching loss, a type of Fisher divergence. In contrast, previous methods often focus on minimizing losses based on KL divergence, consistency, regression, GAN-based adversarial tactics, or a combination thereof.
Appendix C Further Discussion on Long, Short, and Long-Short Guidance
Regarding the relationship between the choice of $\kappa_{2}=\kappa_{3}$ and the categorization into long, short, and long-short guidance, we clarify that various combinations of $\kappa_{1,2,3,4}$ can define these strategies:
Long Strategy: This strategy involves enhancing the teacher’s CFG more than the fake score network during inference or making them equal but both larger than one. It is typically represented by $\kappa_{1}=\kappa_{2}=\kappa_{3}=1,\kappa_{4}>1$, but also includes configurations where $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}>1$.
Short Strategy: This approach aims to reduce the fake-score-network’s CFG during evaluation compared to training or maintain them equal but greater than one. This strategy is exemplified by $\kappa_{1}=\kappa_{4}=1,0<\kappa_{2}=\kappa_{3}<1$, and can also be part of $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}>1$.
Long-Short Strategy: The “simplest” LSG configuration $\kappa_{1}>1,\kappa_{2}=\kappa_{3}=\kappa_{4}=1$ can be interpreted as incorporating both long and short strategies since increasing $\kappa_{1}$ for training the fake score network effectively implies that during inference, the fake score network is guided by a weaker CFG than both the teacher and its own training setting.
SiD-LSG employs a default setting where $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=\kappa>1$. This implies that while the teacher’s training CFG is 1 and evaluation CFG is $\kappa>1$, the fake score network’s training and evaluation CFG are both $\kappa>1$, making the teacher’s CFG during inference stronger than that of the fake score network.
FID-CLIP Compromise: We evaluate various CFG strategies based on the best FID achieved with fewer than 2.56M fake images used to distill generators on Stable Diffusion 1.5, alongside their corresponding CLIP scores. The results, whose trajectories are depicted in Figures 4 and 4 and detailed in Table 3 and Figure 9, indicate that the proposed LSG strategy achieves the best balance between lowering FID and enhancing CLIP scores.
Table 3: Comparison of different CFG strategies in terms of the best FID achieved when iterating with fewer than 2.56M fake images to distill the generators on Stable Diffusion 1.5 and their corresponding CLIP scores.
CFG Strategy $\kappa_{1}$ $\kappa_{2}$ $\kappa_{3}$ $\kappa_{4}$ FID CLIP
Long CFG 1 1 1 2 10.01 0.297
1 1 1 2.5 10.98 0.303
1 1 1 3 11.75 0.303
1 1 1 3.5 12.82 0.304
1 1 1 7.5 15.78 0.308
No CFG 1 1 1 1 15.49 0.269
Short CFG 1 0.5 0.5 1 11.09 0.299
1 0.125 0.125 1 15.54 0.305
“Simplest” LSG 2 1 1 1 11.76 0.306
3 1 1 1 16.38 0.304
LSG 1.5 1.5 1.5 1.5 10.66 0.295
2 2 2 2 10.44 0.304
3 3 3 3 13.88 0.310
4.5 4.5 4.5 4.5 16.69 0.310
Figure 9: Comparison of different CFG strategies in terms of the best FID achieved when iterating with fewer than 2.56M images to distill the generators on Stable Diffusion 1.5 and their corresponding CLIP scores. The markers represent different CFG strategies: Long CFG is denoted by red squares, Short CFG by blue circles, the “simplest” LSG by black diamonds, and LSG by orange stars.
Appendix D Training and evaluation details
@@
The hyperparameters tailored for our study are outlined in Table 4. It’s important to note that the time and memory costs reported in Table 4 do not account for those incurred during periodic evaluations of the FID and CLIP scores of the single-step generator, nor do they include the resources used to save model checkpoints during the distillation process. These costs can vary significantly depending on the computing resources used, including the versions of CUDA and Flash Attention, as well as the storage platforms employed and the frequency of operations.
  Input: Pretrained score network $f_{\phi}$, generator $G_{\theta}$, fake score network $f_{\psi}$, $t_{\text{init}}=625$, $t_{\min}=20$, $t_{\max}=979$, guidance scales $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=1.5$
  Initialization $\theta\leftarrow\phi,\psi\leftarrow\phi$
  repeat
     Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and ${\boldsymbol{c}}$, replacing ${\boldsymbol{c}}$ with the embedding of an empty text string 10% of the time; Define $\boldsymbol{x}_{g}=G_{\theta}(\boldsymbol{z},{\boldsymbol{c}})=f_{\theta}(\sigma_{t_{\text{init}}}\boldsymbol{z},t_{\text{init}},{\boldsymbol{c}})$; Sample $t\in\{t_{min},\ldots,t_{\max}\}$ and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$, and let $\boldsymbol{x}_{t}=a_{t}\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$
  until processing 10M fake images or the training budget is exhausted
  Output: $G_{\theta}$
@@
H100 (80G), FP16 Flash Attention available and enabled Yes
Batch size per GPU 4
# of GPUs 8
# of gradient accumulation round 16
Max memory in GB allocated 32.3 32.2 29.2 35.2
Max memory in GB reserved 32.6 32.4 29.6 35.4
Time in seconds per 1k images 12 12 12 15
Time in hours per 1M images 3 3 3 4
@@
Appendix E Limitations and computational requirements
Memory and speed. Table 4 in Appendix D offers a detailed examination of the computational resources required for the SiD distillation process employing various CFG strategies—Long, Short, “simplest” Long-Short, and the recommended Long-Short—across different NVIDIA GPU platforms (RTX-A5000 with 24GB, RTX-A6000 with 48GB, and H100 with 80GB). Key observations include:
SiD-LSG can be effectively operated on the RTX-A5000, which has 24GB of memory, by enabling xFormers (Lefaudeux et al., 2022) and switching to FP16 for model and gradient precision. On the RTX-A6000 with 48GB of memory, it runs using FP32 with xFormers enabled. However, on our available H100 with 80GB of memory, where xFormers were not supported at the time of our experiments, there was a noticeable increase in GPU memory consumption.
@@ Appendix E Limitations and computational requirements
Further investigation is needed to optimize FP16 performance to match that of FP32. This may involve refining loss scaling techniques, updating packages like Flash Attention or XFormers, or implementing an effective warmup period with FP16 before transitioning to FP32. We plan to address these challenges in our future studies.
We note that our initial experimental platform did not provide proper support for FlashAttention (Dao et al., 2022; Dao, 2023), which adversely affected our FP16 results. Now that we have established proper support for FlashAttention, we are keen to further explore the potential of FP16 in distilling SiD-LSG, especially to determine if it can match FP32’s performance at a lower cost. We have updated Table 4 displayed above to reflect significant memory reductions and a noticeable acceleration in iteration speed under FP16, facilitated by the availability of FlashAttention. This enhancement would enable us to use larger batch sizes per GPU under FP16, further improving time efficiency.
Reaching a performance plateau. Zhou et al. (2024) demonstrate through extensive comparisons that the SiD distilled one-step generator can reduce the FID at an exponential decay rate, showing a log-log linear relationship between the number of iterations and FID. This approach can match or even surpass the performance of both unconditional and label-conditional teacher diffusion models trained under the EDM framework (Karras et al., 2022), provided sufficient training.
However, results shown in Figures 4 and 5 indicate that the SiD-LSG distilled one-step generator on SD1.5 quickly reaches a performance plateau in reducing FID and/or increasing CLIP scores, particularly when the LSG scale exceeds 2. Additionally, data from Table 1 demonstrate that after processing 10M images (approximately 20k iterations at a batch size of 512), SiD-LSG still does not match the text-image alignment performance of the teacher model, which achieves higher CLIP scores after 50 generation steps. Notably, by reducing the LSG scale to 1.5 and doubling the training duration, SiD-LSG achieves a record-low data-free FID of 8.15, establishing a new benchmark among all data-free diffusion distillation methods. This achievement also surpasses the teacher model’s FID of 8.78, which was obtained with a CFG scale of 3 and 50 generation steps. However, this reduction in the LSG scale to 1.5 also results in a significant decline in its CLIP score.
This suggests significant potential for further performance enhancements, possibly by extending beyond single-step generation, increasing the model size, or incorporating real data and additional regression or adversarial losses into the distillation process. These avenues for improvement will be one of the focuses of future studies.
Additional ablation study of the guidance strategies. At the beginning of our research, we evaluated various combinations of $\kappa_{1,2,3,4}$, quickly discarding those that converged too slowly or diverged. Specifically, the setting $\kappa_{1}=1$, $\kappa_{2}=\kappa_{3}>1$, and $\kappa_{4}\in\{\kappa_{2},1\}$ was eliminated early due to its suboptimal performance, as depicted in Figure 11.
Figure 11: This figure shows the progression of FID and CLIP scores during an ablation study of distilling SD1.5 using SiD-LSG, featuring four different sets of $\kappa_{1,2,3,4}$ values. Settings that converged too slowly were terminated early.
Appendix F Prompts and additional example images